# Extended Galerkin neural network approximation of singular variational problems with error control

### v1

@@ 4 Extended Galerkin neural networks with learning of knowledge-based functions
In the numerical examples in Section 3, we approximated the solutions from the set $V_{\mathbf{n},L}^{\sigma}\oplus V_{n_{\lambda}}^{\Phi}$, where $V_{n_{\lambda}}^{\Phi}$ denotes the span of knowledge-based functions $\Phi(x;\lambda)$:
$V_{n_{\lambda}}^{\Phi}:=\{v\;:\;v(x)=\sum_{i=1}^{n_{\lambda}}c_{\lambda,i}\Phi(x;\lambda_{i}),\;\mathbf{c}_{\lambda}\in\mathbb{R}^{n_{\lambda}}\}.$
We utilized exact known values of $\lambda$ when including these functions in the approximation space. However, in many applications some information about the structure of $\Phi$ may be known while the parameters $\lambda$ may be unknown or difficult to compute exactly. In this section, we demonstrate how extended Galerkin neural networks may be applied to the previously studied examples in order to approximate $\lambda$.
@@ 4.1 Poisson equation
$\Phi(r,\theta;\lambda)=r^{\lambda}\sin(\lambda\theta).$
The exact values of $\lambda$ are given by $2/3,4/3,\dots,2j/3,\dots$. We focus only the smallest value of $\lambda$ which describes the asymptotic behavior of $\Phi$ and note that the feedforward part of the neural network, $V_{\mathbf{n},L}^{\sigma}$, is capable of approximating the smoother behavior of $\Phi$ for larger values of $\lambda$.
@@
We again consider the L-shaped domain $\Omega=(-1,1)^{2}\backslash(-1,0)^{2}$ with data $f=1$ and homogeneous Dirichlet boundary conditions. Each Galerkin neural network basis function is obtained according to (23) with $\lambda$ being a trainable parameter. The initial value of $\lambda$ is drawn from the uniform distribution $\mathcal{U}(0,1)$. Figure 14 shows how $\lambda_{1}$ changes value during the backpropagation step of training over 20 independent runs. In each case $\lambda_{1}$ moves toward the true value of $2/3$.
Figure 14: Example 4.1: training of $\lambda_{1}$ and over several runs with initialization $\lambda\sim\mathcal{U}(0,1)$.
4.2 Stokes flow around a nonconvex corner
@@
Figure 15: Eigenvalue training progress for Example 4.2 with initialization $\lambda\sim\mathcal{U}(\lambda_{\text{exact}}-1/3,\lambda_{\text{exact}}+1/3)$.
We approximate the solution to this example using a feedforward network augmented with the knowledge-based function $\Phi(r,\theta;\lambda)=r^{\lambda}\Psi_{\lambda}(\theta)$, where $\Psi_{\lambda}$ is given in Appendix A for the velocity and pressure. We initialize $\lambda$ away from its true value as in Example 4.1 in order to demonstrate that the true eigenvalue of the singular part of the Stokes operator may also be learned by the neural network. Importantly, the singular-behavior of the solution at the origin should be localized with a cutoff function, namely in (21) we take
$\tilde{V}_{m}^{\Phi}:=\left\{v\;:\;v(x)=\sum_{i=1}^{m}c_{\lambda,i}\chi(r)\Phi(x;\lambda_{i}),\;\mathbf{c}_{\lambda}\in\mathbb{R}^{m}\right\},$
where $\chi$ is a $C^{2}$ cutoff function satisfying $\chi(r)=1$ for $r<r_{0}$, $\chi(r)=0$ for $r>r_{1}$.
Figure 15 shows the training progress of $\lambda$ over 30 trials with $\lambda$ initialized according to the uniform distribution $\mathcal{U}(\lambda_{\text{exact}}-1/3,\lambda_{\text{exact}}+1/3)$. We observe that in each trial, $\lambda$ converges to the true value.
@@ 4.3 Stokes flow in a convex corner induced by disturbance at a large distance
We now demonstrate how to learn the knowledge-based functions for the Stokes flow in a convex corner induced by flow at a large distance. The localized domain $\Omega$ is the triangular wedge with vertices $(1,0)$, $(-1,0)$, and $(-3,0)$. The boundary condition for the velocity is a homogeneous Dirichlet condition along the diagonal edges of the wedge and $\mathbf{u}(x,y)=[(1-x)(1+x),0]^{T}$ along the top edge. We again specify $\mathbf{f}=\mathbf{0}$ and $g=0$.
@@
We approximate the solution to this example using a feedforward network augmented with knowledge-based function $\Phi(r,\theta;\lambda)=\mathfrak{Re}[r^{\lambda}\Psi_{\lambda}(\theta)]$. Here, $\lambda$ is assumed to be an unknown complex number of the form $\lambda=\xi+i\zeta$. A cutoff function is again used to isolate the behavior of $\Phi$ to a region around the origin. We initialize $\xi$ and $\zeta$ away from their true values in order to demonstrate that the correct behavior may be learned by the augmented feedforward neural network.
@@ 5 Conclusions
We have presented the extended Galerkin neural network framework for approximating a general class of boundary value problems, including those with low-regularity features. This framework utilizes a simple least squares variational formulation alongside an enriched feedforward neural network architecture. We demonstrate that this approach ensures norm equivalence of the induced energy norm with the norm of the solution space and demonstrate the repercussions when this norm equivalence is lost as is the case when standard least squares approaches based on the $L^{2}$ norm are used. Moreover, we have demonstrated the clear advantage of utilizing knowledge-based functions in the network architecture and have shown that xGNN can easily learn these functions.
@@ A.1 No-slip boundary conditions
$\;\;\;-\cos((\xi_{n}-2)\alpha^{(i)}\cos((\xi_{n}-2)\theta)\cosh(\zeta_{n}\alpha^{(i)})\cosh(\zeta_{n}\theta)$
@@ Appendix B Hyperparameters for numerical examples
Parameter $\mathbf{n}^{(i)}$ $L_{i}$ $\sigma$ learning rate $\delta$
$20\cdot 2^{i-1}$ 1 $\tanh((1+\frac{i}{4})t)$ $0.001/1.1^{i-1}$ $10^{3}$
Table 1: General hyperparameters for learning $\varphi_{i}^{NN}$ in all examples. Any example-specific deviations from this configuration are provided in Tables B.1-B.5.
@@
$n_{\lambda}^{(i)}$222 Only for Example 4.1. $\mathbf{n}^{(i)}/2$
$\Phi$ $r^{\lambda}\sin(\lambda\theta)$
$\lambda$2 $32j/$ 3
@@
$n_{\lambda}^{(i)}$ $5i$ in non-convex corners; $1$ in convex corner
$\Phi$ Equations (A.2) and (50)
@@
of $\Omega$ [^3] The square quadrants are given by $(-2,0)\times(0,2)$ and $(0,2)\times(0,2)$ in the channel. For the triangular cavity, a $128\times 128$ Gauss-Legendre quadrature rule is generated in the rectangle $(-1,1)\times(-3,0)$ and any node falling outside of the cavity has its weight set to 0.; Gauss-Legendre 128 on each edge of $\partial\Omega$
@@
$n_{\lambda}^{(i)}$ 1
$\Phi$ $r^{\lambda}\sin(\lambda\theta)$
@@
$\mathbf{n}^{(i)}$ $n^{(i)}_{\lambda}$ $20\cdot 1.9^{i-1}$ 1
$\Phi$ Equation (48) in 4.2; (45) and (47) in 4.3
@@
tensor-product Gauss-Legendre $128\times 128$ in each square
quadrant of $\Omega$ [^4] In Example 4.3, quadrature rule is generated in the rectangle $(-1,1)\times(-3,0)$ and any points falling outside of $\Omega$ have their weights set to 0.; Gauss-Legendre 128 on each edge of $\partial\Omega$
learning rate $0.003/1.1^{i-1}$

### v2

@@ 4 Extended Galerkin neural networks with learning of knowledge-based functions
In the numerical examples in Section 3, we approximated the solutions from the set $V_{\mathbf{n},L}^{\sigma}\oplus V_{\mathcal{M}}^{\Psi}(\lambda)$ by utilizing exact known values of $\lambda$ when including these functions in the approximation space. However, in many applications some information about the structure of $\Psi$ may be known while the parameters $\lambda$ may be unknown or difficult to compute exactly. In this section, we demonstrate how extended Galerkin neural networks may be applied to the previously studied examples in order to approximate $\lambda$.
@@ 4.1 Poisson equation
$\Psi(r,\theta;\mu)=r^{\mu}\sin(\mu\theta).$
The exact eigenvalues are given by $\lambda_{j}=2j/3$. We focus only on the smallest value of $\lambda$ which describes the asymptotic behavior of $\Psi$ and note that the feedforward part of the neural network, $V_{\mathbf{n},L}^{\sigma}$, is capable of approximating the smoother behavior of $\Phi$ for larger values of $\lambda$.
@@
We again consider the L-shaped domain $\Omega=(-1,1)^{2}\backslash(-1,0)^{2}$ with data $f=1$ and homogeneous Dirichlet boundary conditions. Each Galerkin neural network basis function is obtained according to (29) with $\mu$ being a trainable parameter. The initial value of $\mu$ is drawn from the uniform distribution $\mathcal{U}(0,1)$. Full details of the hyperparameters for this example may be found in Appendix B. Figure 14 shows how $\mu_{1}$ changes value during the backpropagation step of training over 20 independent runs. In each case $\mu_{1}$ moves toward the true value of $2/3$.
Figure 14: Example 4.1: learning of $\lambda_{1}$ and over several runs with initialization $\mu\sim\mathcal{U}(0,1)$.
4.2 Stokes flow around a non-convex corner
@@
Figure 15: Eigenvalue training progress for Example 4.2 with initialization $\mu\sim\mathcal{U}(\lambda_{\text{exact}}-1/3,\lambda_{\text{exact}}+1/3)$.
We approximate the solution to this example using a feedforward network augmented with the knowledge-based function $\Psi(r,\theta;\mu)=r^{\mu}\Psi_{\mu}(\theta)$, where $\Psi_{\mu}$ is given in Appendix A for the velocity and pressure. We initialize $\mu$ away from its true value as in Example 4.1 in order to demonstrate that the true eigenvalue of the singular part of the Stokes operator may also be learned by the neural network. Importantly, the singular-behavior of the solution at the origin should be localized with a cutoff function, namely in (25) we take
$V_{m}^{\Phi}:=\left\{v\;:\;v(x)=\sum_{i=1}^{m}d_{i}\chi(r)\Phi(x;\mu_{i}),\;\mathbf{d}\in\mathbb{R}^{m},\;\lambda_{j}\in\mathcal{I}_{\Phi}\right\},$
where $\chi$ is a $C^{2}$ cutoff function satisfying $\chi(r)=1$ for $r<r_{0}$, $\chi(r)=0$ for $r>r_{1}$. Full details of the hyperparameters for this example may be found in Appendix B.
Figure 15 shows the training progress of $\lambda_{1}$ over 30 trials with $\mu$ initialized according to the uniform distribution $\mathcal{U}(\lambda_{\text{exact}}-1/3,\lambda_{\text{exact}}+1/3)$. We observe that in each trial, $\mu$ converges to the true value.
@@ 4.3 Stokes flow in a convex corner induced by disturbance at a large distance
We now demonstrate how to learn the knowledge-based functions for the Stokes flow in a convex corner induced by flow at a large distance. The localized domain $\Omega$ is the triangular wedge with vertices $(1,0)$, $(-1,0)$, and $(0,-3)$. The boundary condition for the velocity is a homogeneous Dirichlet condition along the diagonal edges of the wedge and $\mathbf{u}(x,y)=[(1-x)(1+x),0]^{T}$ along the top edge. We again specify $\mathbf{f}=\mathbf{0}$ and $g=0$.
@@
We approximate the solution to this example using a feedforward network augmented with knowledge-based function $\Psi(r,\theta;\mu)=\mathfrak{Re}[r^{\mu}\Psi_{\mu}(\theta)]$. Here, $\mu$ is assumed to be an unknown complex number of the form $\mu=\xi+i\zeta$. A cutoff function is again used to isolate the behavior of $\Psi$ to a region around the vertex $(0,-3)$. We initialize $\xi$ and $\zeta$ away from their true values in order to demonstrate that the correct behavior may be learned by the augmented feedforward neural network. Full details of the hyperparameters for this example may be found in Appendix B.
@@ 5 Conclusions
We have presented the extended Galerkin neural network framework for approximating a general class of boundary value problems, including those with low-regularity solutions. The main features of this framework are:
  • The presentation of well-posed least squares variational formulations for arbitrary boundary value problems that are well-suited to approximation by Galerkin neural networks and do not require conversion to first-order systems. We demonstrate that these formulations are dictated by the properties of the PDE under consideration and cannot be chosen arbitrarily by the practitioner.
  • Extended feedforward neural network architectures that systematically incorporate knowledge-based functions into the approximation space and reduce the width of the network required to attain a given accuracy.
  • A rigorous convergence theory which is borne out in the numerical experiments whereas other approaches lack sound theoretical footing and in practice tend to exhibit non-monotonic reduction of the error.
@@ A.1 No-slip boundary conditions
$\;\;\;-\cos((\xi_{n}-2)\alpha^{(i)})\cos((\xi_{n}-2)\theta)\cosh(\zeta_{n}\alpha^{(i)})\cosh(\zeta_{n}\theta)$
@@ Appendix B Hyperparameters for numerical examples
Parameter
$u_{0}(x)$ 0
$\mathbf{n}^{(i)}$ $20\cdot 2^{i-1}$
$L_{i}$ 1
$\sigma_{i}$ $\tanh((1+\frac{i}{4})t)$
@@
Parameter
$u_{0}(x)$ 0
$\mathbf{n}^{(i)}$ $20\cdot 2^{i-1}$
$L_{i}$ 1
$\sigma_{i}$ $\tanh((1+\frac{i}{4})t)$
$\mathcal{M}$111 Only for Example 3.3. 20
$\Psi$ $r^{\mu}\sin(\mu\theta)$
$\mu_{j}$1 $32j/$ 3
@@
Parameter
$u_{0}(x)$ 0
$\mathbf{n}^{(i)}$ $20\cdot 2^{i-1}$
$L_{i}$ 1
$\sigma_{i}$ $\tanh((1+\frac{i}{4})t)$
$\mathcal{M}$ 1 in each non-convex corner and at $(0,-3)$
$\Psi$ Equations (A.2) and (67)
$\mu_{j}$ Equation (59)
@@
of $\Omega$ [^2] The square quadrants are given by $(-2,0)\times(0,2)$ and $(0,2)\times(0,2)$ in the channel. For the triangular cavity, a $128\times 128$ Gauss-Legendre quadrature rule is generated in the rectangle $(-1,1)\times(0,-3)$ and any node falling outside of the cavity has its weight set to 0.; Gauss-Legendre 128 on each edge of $\partial\Omega$
learning rate $3\cdot 10^{-3}/1.1^{i-1}$
$\delta$ $10^{3}$
@@
Parameter
$u_{0}(x)$ 0
$\mathbf{n}^{(i)}$ $40\cdot 2^{i-1}$
$L_{i}$ 1
$\sigma$ $\tanh((1+\frac{i}{4})t)$
$m_{*}$ 1
$\Psi$ $r^{\mu}\sin(\mu\theta)$
$\mu$ Initialized in $\mathcal{U}(0,1)$
@@
Parameter
$u_{0}(x)$ 0
$\mathbf{n}^{(i)}$ $40\cdot 1.9^{i-1}$
$L_{i}$ 1
$\sigma$ $\tanh((1+\frac{i}{4})t)$
$m_{*}$ 1
$\Psi$ Equation (65) in 4.2; (62) and (64) in 4.3
$\mu$
Initialized in $\mathcal{U}(\lambda_{\text{exact}}-1/3,\lambda_{\text{exact}}+1/3)$ for 4.2; initialized
in $\mathcal{U}((\xi_{\text{exact}}-1/2,\xi_{\text{exact}}+1/2)\times(\zeta_{\text{exact}}-1/2,\zeta_{\text{exact}}+1/2))$
for 4.3
@@
tensor-product polar coordinate Gauss-Legendre $128\times 128$ in
$\Omega$ and 512 Gauss-Legendre nodes along $\partial\Omega$ for 4.2;
Gauss-Legendre $128\times 128$ in $\Omega$ [^3] In Example 4.3, quadrature rule is generated in the rectangle $(-1,1)\times(0,-3)$ and any points falling outside of $\Omega$ have their weights set to 0. and Gauss-Legendre 256
on each edge of $\partial\Omega$ in 4.3
learning rate $3\cdot 10^{-3}/1.1^{i-1}$
$\delta$ $10^{3}$